3. The basics of the Capital Asset Pricing Model Aa Aa Which of the following are assumptions of the Capital Asset Pricing Model (CAPM)? Check all that apply. All investors focus on a single holding period All assets are perfectly divisible and liquid. Asset quantities are given and fixed. Standard deviation is the same for all assets Consider the equation for the Capital Asset Pricing Model (CAPM): Cov(ri, rM) 2 In this equation, the term Cov(n, m.) / ơ i represents the covariance between stock i and the market . Suppose that the markets average excess return on stocks is 10.00% and that the risk-free rate is 3.00%. Complete the following table by computing expected returns to stocks for each beta coefficient using the Capital Asset Pricing Model (CAPM): bi -0.50 0.30 1.00 3.00 Return to Stocks (%) -2.00 Based on the CAPM and your calculations for the return to stocks, what does it mean when the coefficient bi = 1? The stock is more volatile than the market. The stock is less volatile than the market. The stocks return correlates with the stock market as a whole.

please answer
Answer #1

1: Assumptions of CAPM are

1, 2 and 3

Investors are looking at a single holding period. Securities can be divided into small portions. They aim to maximize economic utilities -Asset quantities are given and fixed.

2: Cov represents the Covariance between stock and market.

3: Return= Rf+ Beta *Risk premium

I: Return = 3%- 0.5*10% = -2%

II: Return = 3%+ 0.3*10% =6%

III: Return = 3%+1*10% = 13%

IV: Return = 3%+ 3*10% = 33%

4: Stocks return correlates with the stock market as a whole (Since beta is 1)